# Mask Image Watermarking

### v1

@@ 5 Conclusion
We propose MaskMark, a simple, efficient, and flexible framework for image watermarking that supports both global and local watermark embedding, extraction, and localization. It enables applications such as tamper detection and localized image protection. Built on the classical Encoder-Distortion-Decoder paradigm, MaskMark introduces a masking mechanism and a localization module to facilitate accurate and robust watermark extraction, especially in local regions. Extensive experiments demonstrate that MaskMark achieves state-of-the-art performance across all major tasks while maintaining high visual quality and significantly reducing computational cost. We hope MaskMark can serve as a strong foundation and inspire future research in image watermarking.
Appendix A More Details
A.1 Training
@@
A.2 Evaluation
@@
We apply ten valuemetric distortions with the following parameter settings to test generalization:
@@
Geometric Distortions.
We apply three geometric distortions with the following parameter settings to test generalization:
@@
Appendix B More Results
B.1 Visualization Results of Global and Local Watermark Embedding
@@
Figure 11: Visualization results of global watermark embedding using MaskMark-D and local watermark embedding using MaskMark-ED.
B.2 Visualization Results of Localization
@@
Figure 12: More visualization results of watermark localization using different methods.
Figure 13: More visualization results of watermark localization using different methods.
Figure 14: More visualization results of watermark localization using different methods.
Figure 15: More visualization results of watermark localization using different methods.
B.3 Visualization Results of Multiple Watermarks Embedding
@@
B.4 Results of MaskMark-D Adopting Different Visual Quality Enhancement Methods
See Figure 17, Figure 18.
Figure 17: More visualization results of watermark using different visual quality enhancement methods. The residual image is acquired by $10\times|I_{wm}-I_{orig}|$ for observation. Zoom in to see more details.

### v2

@@ 5 Conclusion
In this paper, we propose MaskMark, a simple and efficient framework for both global and local image watermarking. Its core design introduces masks in encoding and decoding stages to guide the encoder and decoder to learn local embedding and extraction. Extensive experiments demonstrate MaskMark ’s superior performance in local watermark extraction and localization, along with high efficiency and adaptability. We hope our simple design can inspire future research to further enhance the practicality and functionality of watermarking models.
Appendix A Limitations
Despite the relatively high PSNR and SSIM of the watermarked images produced by our method, noticeable artifacts are still present in some cases. To address this issue, we explored multiple approaches and ultimately adopted a JND-based method [22] to effectively suppress artifacts (as shown in Appendix D.1). Nevertheless, some images still exhibit visible artifacts after watermark embedding, particularly in smooth background regions. This issue is also present in WAM [19] and reflects a common limitation in local watermarking methods, possibly arising from the inherent trade-off between imperceptibility and local robustness. Our focus in this work is to propose a watermarking framework centered on a mask mechanism, rather than to exhaustively pursue visual quality improvements. We have not further explored advanced strategies such as improved loss functions, auxiliary modules for visual enhancement, or more effective algorithms for watermark strength modulation. These aspects are left for future work to enhance the practicality and visual fidelity of our method.
Appendix B Impact Statement
We propose a simple, efficient, and flexible image watermarking framework whose core component, a masking mechanism, can be easily integrated into existing watermarking models. This mechanism enhances both the functionality and robustness of watermarking systems, significantly improving their practicality in real-world scenarios. The support for local watermark extraction allows reliable recovery of watermark information even when large portions of the image are tampered with. This is particularly valuable for forensic verification in areas such as misinformation mitigation and evidence integrity analysis. The ability to localize watermark regions enables systems to identify precisely which parts of an image contain embedded signals, facilitating transparent content attribution and enabling fair decision-making in legal and copyright-related contexts. Furthermore, local watermark embedding allows protection of specific regions within an image, which is useful when only certain parts require ownership marking or when content from multiple sources needs independent tracking and licensing. By improving spatial controllability while maintaining low computational cost and compatibility with existing architectures, our approach makes robust and fine-grained watermarking more accessible and deployable across diverse applications.
Appendix C More Details
C.1 Training
C.1.1 Details
All images are resized and center-cropped to $256\times 256$ during training. Training is conducted for 100k steps with a batch size of 16 on a single NVIDIA A6000 GPU. We use the AdamW optimizer with a learning rate of $1\times 10^{-4}$, and apply a cosine learning rate scheduler with 2k warm-up steps. We adopt an easy-to-hard training strategy inspired by TrustMark [2]. During the first 0.5k steps, the mask is set to full (i.e., all ones) and no distortion is applied. From step 0.5k to 1k, we introduce all types of masks. After 1k steps, distortions are added. The encoder loss weight $\beta_{\text{enc}}$ is fixed at 1, while the decoder loss weight $\beta_{\text{dec}}$ is initially set to 20 and linearly decayed to 0.2 over the first 5k steps. The mask loss weight $\alpha$ is set to 0.5. The JND module in the encoder is introduced and tuned starting from step 5k, with the scaling factor $\mu$ set to 1.
During fine-tuning with adaptive attacks, VAE-based distortions are applied with a 50% probability, while the original distortion types are retained for the remaining 50%. The hyperparameters are set as follows: $\beta_{\text{dec}}=0.3$, learning rate $=1\times 10^{-4}$, and the quality levels for Bmshj18 and Cheng20 are both set to 5.
C.1.2 Distortions
@@
C.2 Evaluation
C.2.1 Resolution Scaling
Algorithm 1 is adapted from the resolution scaling method described in TrustMark [2]. This algorithm enables a watermark model trained on images with a fixed resolution to be applied at arbitrary resolutions without sacrificing performance.
Input: Original image $\mathbf{x}$, [binary watermark vector $\mathbf{w}$]
Output: Watermarked image $\mathbf{y}$
Data: Embedding network $\mathbf{E}$ trained on the resolution of $m\times n$
1
[…]
$\mathbf{x}\leftarrow\mathbf{x}/127.5-1$ ;
  // Normalize to range [-1, 1]
3 $\bar{\mathbf{x}}\leftarrow\text{interpolate}(\mathbf{x},\ (m,n))$
$\mathbf{r}\leftarrow\mathbf{E}(\bar{\mathbf{x}},\mathbf{w})-\bar{\mathbf{x}}$ ;
  // residual image
4 $\mathbf{r}\leftarrow\text{interpolate}(\mathbf{r},\ (\text{H},\ \text{W}))$
[…]
$\mathbf{y}\leftarrow\mathbf{y}\times 127.5+127.5$
Algorithm 1 Resolution scaling - watermark embedding on arbitrary resolution images
C.2.2 Distortions
@@
We apply ten types of valuemetric distortion with the following parameter settings to evaluate robustness:
@@
Geometric Distortions
We apply three types of geometric distortion with the following parameter settings to evaluate robustness:
@@
C.2.3 Global and Local Watermarking Comparison
Evaluation Setup.
For global watermarking, we sample 1,000 images from the MS-COCO 2014 validation set. We report PSNR and SSIM to assess visual quality, and Bit Accuracy to measure watermark extraction performance. For local watermarking, we construct a comprehensive evaluation set from all 41,000 images in the validation split. We divide the dataset into 12 subsets based on the ratio of masked area to the full image: 1–5%, 5–10%, 10–20%, …, 80–90%, 90–95%, and 95–99%. From each subset, 400 images are randomly selected. To simulate both inpainting and outpainting scenarios, we also include inverted masks, i.e., if the original mask covers $a$$b$of the image, the inverted mask covers $(100{-}b)$$(100{-}a)$%. This yields 800 image-mask pairs per subset and a total of 9,600 for evaluation.
Embedding Strategy.
Baselines and MaskMark-D embed watermarks across the entire image, but the unmasked regions are replaced with the original image to localize the watermark. MaskMark-ED embeds watermark bits only within the masked region and similarly restores the unmasked parts, ensuring true region-specific watermarking.
C.2.4 Multi-Watermark Embedding Setup
Figure 6: Checkerboard-like arrangement of masked regions for multi-watermark embedding. Different colors indicate regions where different watermarks bits are embedded, while black regions contain no embedded information.
Figure 6 illustrates the spatial arrangement of the five non-overlapping masked regions used in our multi-watermark experiments. These regions are fixed at the center, top-left, top-right, bottom-left, and bottom-right of the image, forming a checkerboard-like layout when all five are active. Unlike WAM, which allocates 10% of the image area to each region, we restrict each region to 5%, making the watermark extraction task more challenging. For each configuration (1 to 5 watermarks), we randomly sample 400 images from the MS-COCO 2014 validation set. During evaluation, we use OpenCV’s cv2.connectedComponents to segment the predicted mask into disjoint components, enabling region-wise watermark extraction and scoring.
Appendix D More Results
D.1 Effects of Different Visual Quality Enhancement Methods
Our experiments demonstrate that using only an MSE loss in the pixel space to constrain the difference between the watermarked and original images (referred to as the Base version) already yields high PSNR and SSIM scores. However, despite the favorable metrics, the resulting images often contain visible artifacts that may compromise perceived visual quality. To address this issue, we investigate several enhancement strategies: (1) incorporating a GAN [8] loss, (2) adding a perceptual constraint via LPIPS [27] loss in the feature space, and (3) modulating the watermark signal using a Just Noticeable Difference (JND) [22] module. We conduct experiments on the MaskMark-D variant, but the findings similarly hold for MaskMark-ED.
Figure 7: Visualization results of watermark using different visual quality enhancement methods. Zoom in to see more details. The numbers in parentheses indicate the corresponding loss weights. For example, GAN (0.001) means the GAN loss is assigned a weight of 0.001 in the total loss function, i.e., $\mathcal{L}_{\text{total}}=\beta_{\text{enc}}\mathcal{L}_{\text{enc}}+\beta_{\text{dec}}\mathcal{L}_{\text{dec}}+0.001\mathcal{L}_{\text{GAN}}$, and LPIPS (0.01) indicates that the LPIPS loss is weighted by 0.01, i.e., $\mathcal{L}_{\text{enc}}=\mathcal{L}_{\text{MSE}}(I_{wm},I_{orig})+0.01\mathcal{L}_{\text{LPIPS}}$. To ensure a fair comparison, we adjust either the strength of the added watermark residual or the JND modulation coefficient to maintain comparable PSNR and SSIM across different settings.
As shown in Figure 7, when the weights of the GAN and LPIPS losses are low, they fail to effectively suppress artifacts. Increasing these weights reduces artifacts but adversely affects the performance of local watermark extraction and localization, as illustrated in Figure 8. In contrast, applying JND-based modulation proves more effective: it significantly reduces visible artifacts while maintaining performance comparable to the Base version. These findings suggest that, unlike global perceptual losses such as GAN or LPIPS, JND offers a more adaptive and content-aware modulation strategy. It effectively suppresses visual artifacts while preserving the watermark’s integrity, making it a practical choice for visual quality enhancement.
D.2 Results under Specific Distortions
See Figure 9, Figure 10, Figure 11.
Figure 9: Watermark extraction performance under various specific distortions.
Figure 10: Localization performance of the watermarked area under various specific distortions.
Figure 11: Localization performance of the unwatermarked area under various specific distortions.
D.3 Scalability to Different Watermark Bit Lengths
@@
Figure 12: Watermark extraction performance of MaskMark-D and MaskMark-ED with different bits length. We also show the results of WAM for comparison.
D.4 Enhancing Robustness against Adaptive Attacks via Fast Fine-tuning
See Figure 13.
Figure 13: The effect of VAE fine-tuning on the robustness of MaskMark. Fine-tuning the VAE enhances robustness against VAE attacks, with minimal impact on the original robustness performance.
D.5 Importance of Localization before Extracting Local Watermarks
See Figure 14.
D.6 Visualization Results of Global and Local Watermark Embedding
See Figure 15.
Figure 15: Visualization results of global watermark embedding using MaskMark-D and local watermark embedding using MaskMark-ED. The residual image is acquired by $10\times|I_{wm}-I_{orig}|$ for observation, highlighting the residual more clearly. The same residual visualization strategy is applied in the following figures as well.
D.7 Visualization Results of Localization
@@
Figure 16: Visualization results of watermark localization using different methods. The inpainting results are obtained by applying stable-diffusion-2-inpainting [16] to the masked regions for content reconstruction.
Figure 17: Visualization results of watermark localization using different methods. The inpainting results are obtained by applying stable-diffusion-2-inpainting [16] to the masked regions for content reconstruction.
Figure 18: Visualization results of watermark localization using different methods. The outpainting results are obtained by applying stable-diffusion-2-inpainting [16] to the masked regions for content reconstruction.
Figure 19: Visualization results of watermark localization using different methods. The outpainting results are obtained by applying stable-diffusion-2-inpainting [16] to the masked regions for content reconstruction.
D.8 Visualization Results of Multiple Watermarks Embedding